OpenStudy (anonymous):

What are the values that satisfy the trigonometric equation for 0 < q < 2? \[\sin \theta + \tan( -\theta) = 0\]

Parth (parthkohli):

Use the fact that \(\tan(-\theta) = \tan(\theta)\). You have the following equation:\[\sin(\theta) - \tan(\theta) = 0 \Rightarrow \sin\theta = \tan\theta\]Does that make sense?

OpenStudy (anonymous):

Not really... I barely have any foundation in Trig whatsoever. I'm not sure how to plug anything in..

Parth (parthkohli):

OK, let me explain. Also, I did a typo there: \(\tan(-\theta) = -\tan(\theta) \).

Parth (parthkohli):

First, do you understand the following relation?\[\tan\theta = \dfrac{\sin \theta}{\cos \theta}\]

OpenStudy (anonymous):

To some extent... A friend walked me through this type of problem before.

Parth (parthkohli):

OK. You only need to know that. Also, do you know any trig identities?

OpenStudy (anonymous):

like beta and alpha?

Parth (parthkohli):

Trigonometric identities are some things that are true for all angles. For example, the following...\[\dfrac{\sin\theta}{\cos\theta} = \tan\theta \]...will be true for any \(\theta\) you can imagine. Ever.

OpenStudy (anonymous):

There's other theta's? o.o

Parth (parthkohli):

\(\theta \) is any angle.

Parth (parthkohli):

I'm just saying that an identity is any relation that is true for all angles.

OpenStudy (anonymous):

Oh... got it.